OpenStudy (clarence):

How do you find algebraically the equation of the tangent to y =x^3 + 2x^2 -5x +3 at x = 2?

OpenStudy (anonymous):

Equation of the tangent :D That should already scream "Derivative!!!" But in all seriousness, what's the derivative of this function? ^.^

OpenStudy (clarence):

umm...well, i don't really know :p But i found out that the point of contact is (2,9), is that correct? :p

OpenStudy (anonymous):

Yes. The point of contact is always at (x , f(x) )

OpenStudy (clarence):

So i've done this just faintly, but if the point of contact is (2,9), then dy/dx should be 15?

OpenStudy (anonymous):

yeah, so that's the slope of your tangent line :) So now you have a point (of contact) at (2,9) and a slope m = 15 Can you write the equation of the line now? ^.^

OpenStudy (clarence):

So, is it, 15x-21?? :D

OpenStudy (anonymous):

Yeah, precisely :) y = 15x - 21

OpenStudy (clarence):

WOO!! :D starting to get the hang of this derivatice thingy :) But now it says, "find where the tangent cuts the curve again :O How do you do that?

OpenStudy (anonymous):

Don't know what that means ^

OpenStudy (clarence):

First it says, find the equation of the tangent (which we did) but then it says, find where the tangent cuts the curve again, assuming that curve is the whole y =x^3 + 2x^2 -5x +3 ?

OpenStudy (anonymous):

I'm thinking it's where the tangent line intersects the curve?

OpenStudy (clarence):

how do you find that?

OpenStudy (anonymous):

Set 15x - 21 = x^3 + 2x^2 - 5x + 3

OpenStudy (clarence):

are we suppose to find what x is or?

OpenStudy (anonymous):

Yeah, unfortunately, yes. That's IF I was right in assuming that where the tangent cuts the curve is where it intersects the curve...

OpenStudy (clarence):

well, you haven't been wrong all night yet so :)

OpenStudy (clarence):

but how do you do it? :O I don't know what to do :/

OpenStudy (anonymous):

Me neither :D

OpenStudy (anonymous):

Well, obviously, x = 2 is a solution.

OpenStudy (clarence):

Huh? :p

OpenStudy (anonymous):

Because both the tangent line and the curve are equal to 9, when x = 2 Remember? Because (2,9) is the point of contact.

OpenStudy (clarence):

Yes, so... what now? :p

OpenStudy (anonymous):

Well, rearrange the equation so that 0 is on one side.

OpenStudy (clarence):

and then i have to factorise? :O

OpenStudy (anonymous):

Yes. But don't worry, you already know that one of the factors would be (x-2) so you divide that polynomial by x-2, you will get a quadratic function, which may be factored, or you may use the quadratic formula.

OpenStudy (clarence):

okay, cool thanks :)

OpenStudy (anonymous):

No problem :)